Yomi:

Evaluate: 5n + 3m, when n = 2.5 and m = 1.5: answer choices 17 18 20

Imagine:

So basically solving?

Yomi:

yea ig

Imagine:

So basically it's asking. \[5(2.5)+3(1.5)=?\]

zeze:

14

Yomi:

yes

Yomi:

there is no 14

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zeze 14 \(\color{#0cbb34}{\text{End of Quote}}\) Not 14.

zeze:

3•5 = 15 - 1 = 14

Imagine:

\[12.5+4.5+17.0\]

umm:

\[(5)(2.5)+(3)(1.5)\]\[(5)(2.5)+(3)(1.5)\]\[=17\]

Yomi:

Evaluate: 4.5w - 8.6b, when w = 2 and b = 4 answer choices - 10 - 30.5 - 25.4

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @umm \[(5)(2.5)+(3)(1.5)\]\[(5)(2.5)+(3)(1.5)\]\[=17\] \(\color{#0cbb34}{\text{End of Quote}}\) -High fives-

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yomi Evaluate: 4.5w - 8.6b, when w = 2 and b = 4 answer choices - 10 - 30.5 - 25.4 \(\color{#0cbb34}{\text{End of Quote}}\) Please ask it on a different question.

umm:

\[(4.5)(2)−(8.6)(4)\]\[(4.5)(2)−(8.6)(4)\]\[=−25.4\]

Yomi:

i have like 7 of them i need to ask and i cant keep doing like more than 3 for the same subject